OpenStudy (anonymous):

how do I know if (-3,4) is a solution to 5x-2y=-23? please show me how to work this, thanks.

OpenStudy (anonymous):

Put x = -3 in the equation & see if y turns out to be 4. :)

OpenStudy (anonymous):

Yes it does turns out to be 4. -3,4 is a solution of the given equation. :D

OpenStudy (anonymous):

Hey there, (-3,4) is in the form of (x,y). So 'x=-3' and 'y=4'. Simply plug these values into the equation and see if the left side equals the right side. If it does, then it is a solution.

OpenStudy (anonymous):

5(-3) -2y = -23 2y = -15 - (-23) 2y=8 y=4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

put (-3,4) in ur equation if u get 0 this mean (-3,4) is a soln.